anysize.sty — to set the margin sizes in L^{A}TEX

first version: Michael Salzenberg, conversion to L^{A}TEX2e: Thomas Esser, documentation: Dirk Hillbrecht^{∗}

v1.0, Aug 13, 1994

1 Overview

anysize is a rather small style for L^{A}TEX to adjust the sizes of the paper and the margins of a document. These settings are done through one simple command.

anysize is full L^{A}TEX2e–compilant and uses the new features of this L^{A}TEX–version. Therefore, it will not run under L^{A}TEX 2.09.

Many thanks go to Michael Salzenberg, who made the first implementation of this style. He does not want to be bothered with the wide-area-publication of the style, so Thomas — who converted it to L^{A}TEX2e — and me, who documented it properly (I hope :-)), maintain it now. Therefore there are only our e-Mail-adresses given.

2 Usage

anysize is included through the \usepackage–directive. The style knows about the options on page size (a4paper, a5paper etc.) and a further option germanpar, that changes the margins of the paragraphes in the document particulary.

After inclusion of anysize, there are two new commands available for the document preamble:

• \papersize{width}{height}: sets the paper size. This command should not be used anymore, as there are standardized style options for the standard paper sizes to be passed through the \documentclass–command in the header of the document.

• \marginsize{left}{right}{top}{bottom}: sets the margin sizes. This is the really interesting thing of the whole one. In two–sided documents, left and right margin are those ones for the rectos. For the backs, they are swapped.

These are the default settings, as long as none of the above commands is given:

paper size^{1}:

width 21,0 cm
height 29,7 cm

(DIN A4) margin sizes:

left  3,0 cm
right  2,0 cm
top  2,0 cm
bottom 2,0 cm

∗eMail–addresses: te@informatik.uni-hannover.de, hillbrec@informatik.uni-hannover.de

1only if there is no style option for the paper size